OpenStudy (manymath09):

Jenna has a bag that contains 7 red marbles and 25 blue marbles. She selects a marble at random, and then, without replacing the first one, selects another marble at random. What is the probability that Jenna selects a blue marble and then a red marble? Round your answer to the nearest percent. P(blue and red) ≈

OpenStudy (.mehek.):

Blue: \(\dfrac{25}{32}\) When you don't replace, the total goes down by 1 Red: \(\dfrac{7}{31}\) Multiply both \(\dfrac{25}{32}*\dfrac{7}{31}=\)

OpenStudy (manymath09):

175/992

OpenStudy (manymath09):

?

OpenStudy (.mehek.):

Use a calculator to divide and multiply by 100

OpenStudy (manymath09):

idk please help

OpenStudy (.mehek.):

Do you have a calculator?

OpenStudy (manymath09):

no

OpenStudy (.mehek.):

On your computer?

OpenStudy (manymath09):

no it does not work

OpenStudy (manymath09):

ok what do I do

OpenStudy (.mehek.):

put 175/992

OpenStudy (manymath09):

ok

OpenStudy (.mehek.):

and then multiply by 100

OpenStudy (.mehek.):

Tell me what you get

OpenStudy (manymath09):

17.6411290323

OpenStudy (.mehek.):

round to the nearest percent

OpenStudy (manymath09):

17.6?

OpenStudy (.mehek.):

nearest percent not tenth

OpenStudy (manymath09):

18?

OpenStudy (.mehek.):

Yes

OpenStudy (manymath09):

thank you sooo much

OpenStudy (.mehek.):

You're welcome :)